# Fragile Preferences: A Deep Dive into Order Effects in Large Language Models

### v1

@@ 2.4 Interaction of Positional Bias with Other Biases
The positional bias in triplewise color comparisons is so strong that in every tier, there is at least one position that is never chosen at $T=0$ (except in GPT-4o-mini’s ‘Fair’ tier, which selected the last position a single time). This positional skew appears less pronounced in resume comparisons—particularly for Claude 3 Haiku and Llama 3 8B. Upon closer inspection, this reduction in positional bias appears to be driven by strong, competing biases for or against specific synthetic personas, especially individual names. Claude 3 Haiku, in particular, exhibits strong name-based preferences: it consistently favors certain names over others, regardless of resume content or position. For example, in direct comparisons, Claude 3 Haiku ‘Christopher Taylor’ over ‘Andrew Harris’ in 64% of the cases (82 out of 128 trials, binomial $p=0.0019,h=0.29$). These patterns emerged even though all names were drawn from a controlled set of generic Caucasian U.S. identities, chosen to minimize racial, ethnic, and other socially marked biases that have been documented in LLMs [41, 42].
This pattern is evident in 3(a), which shows the distribution of selection frequencies for each name across triplewise comparisons. Claude 3 Haiku’s distribution is flat, indicating that a few names are chosen disproportionately often while others are rarely selected—consistent with strong and persistent name preferences. GPT-4o-mini, by contrast, displays a near-perfect normal distribution, suggesting minimal name-based bias in the triplewise setting. Llama 3 8B lies between these extremes, with moderate variation across names. Although GPT-4o-mini shows no signs of name bias in triplewise comparisons, we detect a small but statistically significant name preference in the pairwise setting. See Appendix H for full details.
@@
Prior work on gender bias in algorithmic hiring is mixed: some studies report little to no bias [43], others find systematic preferences for male candidates [44, 7], while more recent evaluations suggest a tilt toward female candidates [8, 9] or context-dependent outcomes [13]. Studies using pairwise evaluations have found a bias favoring female candidates [35, 8, 9], a pattern we also replicate: for example, GPT-4o-mini selects female candidates 55.7% of the time at $T=1$ ($p=3.22\times 10^{-4}$, $h=0.11$), with Claude 3 Haiku and Llama 3 8B showing similar trends.
While gender bias in algorithmic decision-making is an important topic in its own right, our primary goal is to examine how gender interacts with positional order effects. We find that gender effects are overshadowed by substantially stronger order biases. Across all three models, the candidate in the dominant position (first or second, as defined by same-gender comparisons) is selected at much higher rates. For GPT-4o-mini, this rate reaches 92.8% in mixed-gender comparisons—compared to 93.6% in same-gender trials—suggesting that gender bias can at times counteract, but rarely override, the positional preference. Claude 3 Haiku and Llama 3 8B show similar patterns, with positional bias rates of 84.3% and 72.3% in mixed-gender comparisons, versus 87.6% and 89.2% in same-gender comparisons, respectively. Bootstrapped confidence intervals confirm that order bias significantly exceeds gender bias across all models (95% CI for $h_{\text{order}}-h_{\text{gender}}$ excludes 0 in all cases).
This difference between the effect sizes of the positional and gender effects can be clearly seen in 3(b), where gender effects manifest in two ways: a modest overall preference for female candidates, and the observation that nearly all instances of overcoming the positional disadvantage involve female candidates. Full statistics appear in Appendix I.
@@ 3 Discussion
This study demonstrates systematic positional biases in LLM-driven comparisons, with methodological, theoretical, and practical implications. Across both color and hiring domains, we find that positional biases are quality-dependent and strong: primacy effects dominate for high-quality options, while lower-quality choices elicit a recency effect.
Most prior research has not clearly distinguished between two possible sources of positional effects: arbitrary tie-breaking and genuine preference distortion. Many studies have observed that order effects appear when the model is indifferent between the different options [31, 17, 32], and studies on positional biases and the decoy effect often purposefully use alternatives of similar quality  [45, 37]. While a few human studies have attempted to disentangle these explanations [46], conclusions remain limited due to humans’ limited ability to determine their ‘true’ preferences [24] or compute expected utilities [47]. The difficulty to determine ground-truth preferences in human subjects is further complicated by the fact that preferences are influenced by contextual factors and shift over time [48].
@@
Much of the work on LLM bias has centered on uncovering and correcting cognitive distortions, such as anchoring, confirmation bias, or gender stereotyping, that models acquire from human-generated training data [49, 50, 51, 41]. These studies often argue that models inherit cognitive biases that are present in the data [52, 53, 54]. Even studies that identify behavior that diverges from human norms (for instance, a female‐candidate preference in resume‐ranking tasks, where humans often favor male candidates) still frame these findings through the lens of human cognitive biases [8, 9]. In contrast, we uncover two new failure modes that have no clear analogue in human decision-making: a centrality bias in triplewise comparisons and a name bias, even though names were explicitly chosen to avoid known demographic triggers.
@@
These patterns raise new theoretical questions. Classical explanations for human order effects, such as memory limitations [55, 56], proactive inhibition [57], or heuristics [24], cannot readily explain our findings of quality-dependent bias shifts or the centrality bias. One possible account—drawing on the long-established notions of reference points [40, 58] and contrast effect [26, 27] in behavioral decision theory—is that LLMs establish an implicit reference point from their training data and then evaluate each option in contrast to this baseline. Above-baseline options are inflated, leading to primacy, while below-baseline ones are penalized. However, it is not clear how to extend this explanation to biases observed in triplewise comparisons, in particular the centrality bias. Moreover, model-specific differences (e.g., centrality effects in GPT-4o-mini and Claude 3 Haiku but not Llama 3 8B), indicate that different architectures details may give rise to distinct bias mechanisms. Together, these findings call for new theoretical frameworks to explain how biases emerge and propagate in artificial agents—frameworks that go beyond human cognitive models and the biases embedded in training data.
From a practical standpoint, standard mitigation strategies such as randomizing orders or averaging across queries do not fully eliminate positional distortions. Instead, we propose approaches that leverage LLM-specific characteristics. First, raising the temperature uncovers fragile preferences that deterministic (zero-temperature) settings suppress. Although low temperatures are often favored for precision and factuality, and high temperatures for creativity [38, 39], we find that in these comparison tasks a higher temperature can actually improve selection accuracy. Second, positional bias itself can be informative: a pronounced primacy effect may indicate that both options are of high quality, whereas a strong recency effect may signal that both are of lower quality.
@@
While our findings hold across three contemporary models, new architectures should be audited for emergent order and identity biases. In high-stakes domains such as legal reasoning, medical diagnosis, or scientific peer review, bias-aware evaluation pipelines can combine model-driven diagnostics with human-in-the-loop checks. Finally, causal interpretability methods and adversarial probing could illuminate the internal mechanisms that give rise to preference distortions, paving the way toward bias-robust LLM design and deployment.
@@ 4.1 Resume and Color Generation and Selection
As order effects are more pronounced when the options are of comparable quality, we generated three options for each of four quality tiers of colors and resumes. For colors, we denoted these tiers by ‘Ideal’, ‘Fair’, ‘Plain’ and ‘Harsh’; for resumes, by ‘Best’, ‘Good’, ‘Mediocre’ and ‘Weak’. We chose different names for the tiers in each domain so as not to imply an equivalence between them (i.e., ‘Plain’ colors are not necessarily comparable to ‘Mediocre’ resumes in terms of quality).
We first generated the candidates for each tier. For the resumes, we had GPT-4o-mini generate nine sets of four resumes of decreasing quality for each of four professions: Mechanical Engineer, Registered Nurse, Journalist, and Real Estate Agent. We removed the personal information from the resumes. The resume generation prompts are shown in Appendix B. For colors, we asked each model to generate colors that they ‘thought’ were comparable.
For each quality tier, we performed triplewise comparisons among the candidate options and selected three options based on the following criteria. First, each had to be selected in at least one of the six possible permutations. Second, to ensure that the tier labels were meaningful, we checked that colors from a higher tier were chosen over those from a lower tier in at least two-thirds of all pairwise comparisons. To select the colors for pairwise comparisons in each tier, we evaluated all three possible pairs among the three options and selected a pair in which neither color was consistently preferred over the other. We note that our selection methodology led to the fact that we did not necessarily use the same resumes or colors for the three models.
For the personal information, We generated 30 fictional personas, using 15 common American Caucasian last names, male first names and female first names, selected from the list of the most popular names in the US as tracked by the SSA [59]. We paired each first name with a random surname and state, and randomly extrapolated to a complete profile. For more details, see Appendix B.
@@ 4.2 Pairwise Comparisons
For each pair, we presented the model with both orders to isolate position effects. For resumes, we selected four male names and four female names and performed exhaustive same-gender comparisons. That is, for each pair of names $i$ and $j$, we compared resume 1 with $i$’s personal information to resume 2 with $j$’s information, and vice versa. This yielded a total of 768 pairwise comparisons per model at each temperature ($4$ professions $\times$ $4$ tiers $\times$ $2$ genders $\times$ $\binom{4}{2}=6$ name pairs $\times$ $2$ name-resume assignments $\times$ $2$ presentation orders). For colors, we repeated the prompt for each permutation 50 times at each temperature $T\in\{0,0.5,1\}$.
For the cross-gender comparisons, we used the same resumes and personal information as in the same-gender comparisons and compared all combinations of male and female names. That is, for each male name $i$ and female name $j$, we compared resume 1 with $i$’s personal information to resume 2 with $j$’s information, and vice versa. This yielded 1,024 additional pairwise comparisons per model at each temperature ($4$ professions $\times$ $4$ tiers $\times$ $4$ female names $\times$ $4$ male names $\times$ $2$ name-resume assignments $\times$ $2$ presentation orders).
@@ 4.3 Triplewise Comparisons
For each resume triple, we randomly assigned the 15 male and 15 female names (and their associated personal information) across the three resumes, yielding 5 distinct triples per gender. We then presented all 6 permutations of each triple to each LLM, asking the model to select the strongest candidate; the exact prompt is provided in Appendix B. In total, we conducted 960 triplewise comparisons per model at each temperature ($4$ professions $\times$ $4$ tiers $\times$ $2$ genders $\times$ $6$ permutations $\times$ $5$ sets of triples). For colors, we repeated the prompt for each permutation 40 times at each temperature, for a total of 240 prompts. the exact prompt for color comparisons is provided in Appendix A.
4.4 Data Cleaning
Although we instructed the model to output only the name of the strongest candidate in each comparison, LLMs occasionally returned more verbose or inconsistent responses (e.g., “Candidate 1: Amanda Thomas”). In both the color and resume domains, we corrected clearly identifiable output errors and discarded nonsensical responses (e.g., Claude 3 Haiku selecting “White” in a comparison between Burgundy and Hot Pink). However, we did not discard all mismatches, as some models—Claude 3 Haiku in particular, but also Llama 3 8B—occasionally produced minor spelling variations (e.g., “Offwhite” instead of “Off-White,” or “Palelinen” instead of “Pale Linen”). Discarding all such cases would introduce bias; for example, if Off-White was misspelled while Beige was not, discarding those comparisons would unfairly favor Beige.
@@ 4.5 Statistical Tests
We evaluated positional order effects in pairwise comparisons using two-tailed binomial tests, assessing whether the proportion of selections for the first versus second position significantly deviated from the null hypothesis of no positional bias (equivalent to $p=0.5$). As all other effects are controlled for, a statistically significant deviation indicates that the model’s choices are systematically influenced by presentation order. We report effect sizes using Cohen’s $h$, a standardized measure of deviation from the expected proportion.
For triplewise comparisons, we used chi-square goodness-of-fit tests to evaluate whether the distribution of selections across the three positions departed from a uniform distribution (i.e., each position selected with probability $1/3$ under the null hypothesis of no bias). Significant deviations indicate systematic positional preferences. Effect sizes for these tests are reported using Cramér’s $V$. We did not perform statistical tests for comparisons at $T=0$ as the results are almost deterministic.
To compare the relative strength of gender and positional biases, we used the same dataset and performed nonparametric bootstrapping. For each of 10,000 resamples, we computed effect sizes for both gender and positional bias within the same matched examples. The resulting bootstrap distribution allowed us to assess the consistency and magnitude of the difference, confirming that order effects consistently exceeded gender effects across resamples.
4.6 Data, Materials, and Software Availability
Replication archive with code and data is available at Open Science Framework at https://osf.io/59df6/.

### v2

@@ 2.4 Interaction of Position Bias with Other Biases
The position bias in triplewise color comparisons is so strong that in every tier (2(a)), at least one position was never chosen at $T=0$ (except in GPT-4o-mini’s Fair tier, which selected the last position a single time). This positional skew was less pronounced in resume comparisons, particularly for Claude 3 Haiku and Llama 3 8B. Upon closer examination, this reduction in position bias appears to have been driven by strong, competing biases for or against specific synthetic personas, especially individual names. Claude 3 Haiku, in particular, exhibited strong name-based preferences: it consistently favored certain names over others, regardless of resume content or position. For example, in direct comparisons, Claude 3 Haiku selected ‘Christopher Taylor’ over ‘Andrew Harris’ in 64% of the cases (82 out of 128 trials, binomial $p=0.0019,h=0.29$). These patterns emerged even though all names were drawn from a controlled set of generic Caucasian U.S. identities, chosen to minimize racial, ethnic, and other socially marked biases that have been documented in LLMs [44, 45].
This pattern is evident in 3(a), which shows the distribution of selection frequencies for each name across triplewise comparisons. Claude 3 Haiku’s distribution is flat, indicating that a few names were chosen disproportionately often while others were rarely selected—consistent with strong and persistent name preferences. GPT-4o-mini, by contrast, displays a near-perfect normal distribution, suggesting minimal name-based bias in the triplewise setting. Llama 3 8B lies between these extremes, with moderate variation across names. Although GPT-4o-mini showed no signs of name bias in triplewise comparisons, we detected a small but statistically significant name preference in the pairwise setting. See Appendix I for full details.
@@
Prior work on gender bias in algorithmic hiring is mixed: some studies report little to no bias [46], others find systematic preferences for male candidates [47, 7], while more recent evaluations suggest a tilt toward female candidates [8, 9] or context-dependent outcomes [13]. Studies using pairwise evaluations have found a bias favoring female candidates [35, 8, 9], a pattern we also replicated: for example, GPT-4o-mini selected female candidates 54.8% of the time at $T=0$, with Claude 3 Haiku and Llama 3 8B showing similar trends.
While gender bias in algorithmic decision-making is an important topic in its own right, our primary goal is to examine how gender interacts with positional order effects. We found that gender effects were overshadowed by substantially stronger order biases. Across all three models, the candidate in the dominant position (first or second, as defined by same-gender comparisons) was selected at much higher rates. For GPT-4o-mini, this rate reached 92.8% in mixed-gender comparisons—compared to 93.6% in same-gender trials—suggesting that gender bias can at times counteract, but rarely override, the positional preference. Claude 3 Haiku and Llama 3 8B showed similar patterns, with position bias rates of 84.3% and 72.3% in mixed-gender comparisons, versus 87.6% and 89.2% in same-gender comparisons, respectively. Bootstrapped confidence intervals confirm that order bias significantly exceeds gender bias across all models (95% CI for $h_{\text{order}}-h_{\text{gender}}$ excludes 0 in all cases); see SI Appendix J for figures and additional details.
This difference between the effect sizes of the positional and gender effects can be clearly seen in 3(b), where gender effects manifest in two ways: a modest overall preference for female candidates, and the observation that nearly all instances of overcoming the positional disadvantage involve female candidates. Full statistics appear in Appendix J.
@@ 3 Discussion
This study provides evidence of systematic position biases in LLM-driven comparisons, with methodological, theoretical, and practical implications. Across both color and hiring domains, we find that position biases are quality-dependent and strong: primacy effects dominate for high-quality options, while later options (second or third) are preferred for lower-quality choices.
Most prior research has not clearly distinguished between two possible sources of positional effects: arbitrary tie-breaking and genuine preference distortion. Many studies have observed that order effects appear when the model is indifferent between the different options [31, 17, 32], and studies on position biases and the decoy effect often purposefully use alternatives of similar quality [48, 37]. While some human studies have attempted to disentangle these explanations [49], conclusions remain limited due to humans’ limited ability to determine their ‘true’ preferences [24] or compute expected utilities [50]. Determining ground-truth preferences in human subjects is further complicated by contextual influences and shifts in preferences over time [51].
@@
Much of the work on LLM bias has centered on uncovering and correcting cognitive distortions, such as anchoring, confirmation bias, or gender stereotyping, that models acquire from human-generated training data [52, 53, 54, 44]. These studies often argue that models inherit cognitive biases that present in the data [55, 56, 57]. Even when studies identify behavior that diverges from human norms (for instance, a female‐candidate preference in resume‐ranking tasks, where humans often favor male candidates), these findings are often interpreted through the lens of human cognitive biases [8, 9]. In contrast, we uncover two new failure modes that have no clear analogue in human decision-making: a centrality bias in triplewise comparisons and a name bias, even though the names were explicitly chosen to avoid known demographic associations.
@@
These patterns raise new theoretical questions. Classical explanations for human order effects, such as memory limitations [58, 59], proactive inhibition [60], or heuristics [24], cannot readily explain our findings of quality-dependent bias shifts or the centrality bias. One possible account—drawing on the long-established notions of reference points [42, 61] and contrast effect [26, 27] in behavioral decision theory—is that LLMs establish an implicit reference point from their training data and then evaluate each option in contrast to this baseline. Above-baseline options are inflated, leading to primacy, while below-baseline ones are penalized. However, it is not clear how to extend this explanation to biases observed in triplewise comparisons, in particular the centrality bias. Moreover, model-specific differences (e.g., centrality effects in GPT-4o-mini and Claude 3 Haiku but not Llama 3 8B) suggest that distinct bias mechanisms may arise across models, potentially due to differences in architecture or training, though the precise sources of these differences, especially in closed models, remain unclear. Together, these findings call for new theoretical frameworks that go beyond human cognitive models in explaining how biases emerge and propagate in artificial agents.
Our findings also carry practical implications. It is critical to not only monitor for human-like cognitive distortions inherited through training data, but also to proactively identify and mitigate novel failure modes that emerge from the models’ architectures or training dynamics. These uniquely artificial biases may be harder to anticipate precisely because they have no direct human analogue. This underscores the importance of robust evaluation practices tailored to large language models.
We find that position biases are generally stronger than gender biases. This might be due to alignment efforts that explicitly target and reduce social biases like gender, while largely overlooking position biases. Nevertheless, because positional effects currently overshadow gender-based biases, their mitigation is equally, if not more, critical.
However, standard mitigation strategies such as randomizing orders or averaging across queries do not fully eliminate order effects: randomizing the order simply randomizes the option that benefits from the bias, and averaging often assigns all options equal scores when the bias is strong. [^5] For instance, in pairwise comparisons, both options will be chosen an equal number of times when the first option is always chosen. To address these effects, we propose two LLM-specific mitigation approaches:
  1. 1.
    Perform repeated queries at a higher temperature to uncover suppressed preferences. Due to near-determinism at $T=0$, position bias can completely mask fragile preferences, rendering them indistinguishable from true indifference. Increasing the temperature introduces controlled stochasticity that allows latent preferences to emerge and improves selection accuracy.
    While we only partitioned strict preferences into robust and fragile, repeated querying allows us to quantify the strength of the preference. If a model prefers option $a$ over option $b$ 60% of the time at $T=1$, and option $c$ over option $d$ 80% of the time, we can infer that $a\succ b$, $c\succ d$ and furthermore that $c\succ d$ is a stronger preference than $a\succ b$. Repeated testing allows us to separate these cases with high statistical confidence.
    While low temperatures are often favored for precision and factuality and higher temperatures are reserved for creative generation [39, 40], our findings show that in comparison tasks, higher temperatures can enhance precision by revealing the model’s underlying preferences.
  2. 2.
    Use position bias as a diagnostic signal. When a strong primacy effect is detected, it may signal that both options are of high quality; conversely, a strong recency effect may suggest that both are of lower quality. This heuristic can guide downstream decision rules, for instance, by triggering additional evaluation or human review when strong order effects are observed.
These strategies treat position bias not only as a problem to be corrected but also as a source of information. In practice, they can be incorporated into evaluation pipelines by running targeted high-temperature queries to detect fragile preferences, and by interpreting the magnitude and direction of position bias as indicators that guide further decision-making. Crucially, our experimental results show that these approaches can recover suppressed preferences and reduce the likelihood of selecting an inferior option due solely to presentation order.
@@
Our results hold across two disparate domains, suggesting that they are inherent rather than domain-specific. Nevertheless, new architectures should be audited for order and identity biases. In high-stakes domains such as legal reasoning, medical diagnosis, or scientific peer review, bias-aware evaluation pipelines can combine model-driven diagnostics with human-in-the-loop checks. Finally, causal interpretability methods and adversarial probing could illuminate the internal mechanisms that give rise to preference distortions, paving the way toward bias-robust LLM design and deployment.
@@ 4.1 Resume and Color Generation and Selection
As order effects are more pronounced when the options are of comparable quality, we generated three options for each of four quality tiers of colors and resumes. For colors, we denoted these tiers by Ideal, Fair, Plain, and Harsh; for resumes, by Best, Good, Mediocre, and Weak. We chose different names for the tiers in each domain so as not to imply an equivalence between them (i.e., Plain colors are not necessarily comparable to Mediocre resumes in terms of quality).
We first generated the candidates for each tier. For the resumes, we had GPT-4o-mini generate nine sets of four resumes of decreasing quality for each of four professions: Mechanical Engineer, Registered Nurse, Journalist, and Real Estate Agent. We removed the personal information from the resumes. For colors, we asked each model to generate lists of colors that they ‘thought’ were comparable.
To validate the tier labels, we compared all items in each tier with all items in the next lower tier (highest vs. second-highest, second-highest vs. third-highest, and third-highest vs. lowest). We verified that, in each case, the higher-tier colors were more likely to be selected (one-sided binomial test, null hypothesis: $p=0.5$) with $p$-values of at most $0.001$.
For each quality tier, we performed triplewise comparisons among the candidate options and selected three options such that each was selected in at least one of the six possible permutations. To select the colors for pairwise comparisons in each tier, we evaluated all three possible pairs among the three options and selected a pair in which neither color was consistently preferred over the other. We note that our selection methodology led to the fact that we did not necessarily use the same resumes or colors for the three models.
For the personal information, We generated 30 fictional personas, using 15 common American Caucasian last names, male first names and female first names, selected from the list of the most popular names in the US as tracked by the SSA [62]. We paired each first name with a random surname and state, and randomly extrapolated to a complete profile.
@@ 4.2 Pairwise Comparisons
For each pair, we presented the model with both orders to isolate position effects. For resumes, we selected four male names and four female names and performed exhaustive same-gender comparisons. That is, for each pair of names $i$ and $j$, we compared resume 1 with $i$’s personal information to resume 2 with $j$’s information, and vice versa. This yielded a total of 768 pairwise comparisons per model at each temperature ($4$ professions $\times$ 44 tiers $\times$ 22 genders $\times$ $\binom{4}{2}=6$ name pairs $\times$ 22 name-resume assignments $\times$ 22 presentation orders). For colors, we repeated the prompt for each permutation 50 times at each temperature $T\in\{0,0.5,1\}$.
For the cross-gender comparisons, we used the same resumes and personal information as the same-gender comparisons and compared all combinations of male and female names. That is, for each male name $i$ and female name $j$, we compared resume 1 with $i$’s personal information to resume 2 with $j$’s information, and vice versa. This yielded 1,024 additional pairwise comparisons per model at each temperature ($4$ professions $\times$ 44 tiers $\times$ 44 female names $\times$ 44 male names $\times$ 22 name-resume assignments $\times$ 22 presentation orders).
@@ 4.3 Triplewise Comparisons
For each resume triple, we randomly assigned the 15 male and 15 female names (and their associated personal information) to the three resumes, yielding 5 distinct triples per gender. We then presented all 6 permutations of each triple to each LLM, asking the model to select the strongest candidate. In total, we conducted 960 triplewise comparisons per model at each temperature ($4$ professions $\times$ 44 tiers $\times$ 22 genders $\times$ 66 permutations $\times$ 55 sets of triples). For colors, we repeated the prompt for each permutation 40 times at each temperature, for a total of 240 prompts.
@@ 4.4 Statistical Tests
We evaluated positional order effects in pairwise comparisons using two-tailed binomial tests, assessing whether the proportion of selections for the first versus second position significantly deviated from the null hypothesis of no position bias (equivalent to $p=0.5$). As all other effects are controlled for, a statistically significant deviation indicates that the model’s choices were systematically influenced by presentation order. We report effect sizes using Cohen’s $h$, a standardized measure of deviation from the expected proportion.
For triplewise comparisons, we used chi-square goodness-of-fit tests to evaluate whether the distribution of selections across the three positions departed from a uniform distribution (i.e., each position selected with probability one third under the null hypothesis of no bias). Significant deviations indicate systematic positional preferences. Effect sizes for these tests are reported using Cramér’s $V$. We did not perform statistical tests for comparisons at $T=0$ as the results are almost deterministic.
To compare the relative strength of gender and position biases, we used the same dataset and performed nonparametric bootstrapping. For each of 10,000 resamples, we computed effect sizes for both gender and position bias within the same matched examples. The resulting bootstrap distribution allowed us to assess the consistency and magnitude of the difference, confirming that order effects consistently exceeded gender effects across resamples.